OpenStudy (anonymous):

Trinh is catering a wedding that 112 people will attend. She is making two cakes, a chocolate cake and a carrot cake. 5/8of the guests will be served chocolate cake and the rest will receive carrot cake. How many guests will get carrot cake?

OpenStudy (mertsj):

If 5/8 get chocolate, then 3/8 get carrot. So can you find 3/8 of 112?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Create a proportion.

OpenStudy (mertsj):

In mathematics, "of' means times. so 3/8 x 112

OpenStudy (anonymous):

lol yea i was gonna suggest: 112 - 1 x - 3/8; so x = (112*(3/8))/1 = 112*(3/8)

OpenStudy (anonymous):

42

OpenStudy (mertsj):

good for you.

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw